OpenStudy (anonymous):

HELP PLEASE. Solve using two different methods. Explain which method you found to be more efficient. y=1/2x-6 2x + 6y=19 I have no idea.. ):

OpenStudy (anonymous):

@Luigi0210 can you help me?

OpenStudy (anonymous):

y=-1/2 x=11. Sub it into 2x+6y=19 to check it.

OpenStudy (anonymous):

Omg thank you!

OpenStudy (anonymous):

no prob. close the question